OpenStudy (anonymous):

0.57142857142 as a fraction

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

if you look at this number the repeatition of digits start after 2 now you have to move the decimal point exactly to the right of digit 2

OpenStudy (anonymous):

let x = 0.57142857142 multiply x by 1000000 we get 1000000x = 571428.571428 1000000x - x = 571428 999999x = 571428 x = 571428/999999

OpenStudy (anonymous):

The answer above can be simplified as \[ \frac{571428} {999999}=\frac 4 7 \]